OpenStudy (anonymous):

Simplify: -6(7+x) -3(x-5) -5(10x+1) (1+2v)x 5

OpenStudy (anonymous):

-42 - 6x -3x + 15 -50x + 5 5 + 10v

OpenStudy (anonymous):

are the answers when multiplied ;)

OpenStudy (anonymous):

*-50x -5

OpenStudy (anonymous):

-6(7+x) = -42 - 6x -3(x-5) = -3x + 15 -5(10x+1) = -50x - 5 (1+2v)x 5 = 5 + 10v